OpenStudy (anonymous):

Solve 6 over x minus 4 equals 4 over x x for x and determine if the solution is extraneous or not. x = -8, extraneous x = -8, non-extraneous x = 8, extraneous x = 8, non-extraneous

OpenStudy (jack1):

6 over x minus 4 equals 4 over x x \[\Large \frac {6}{x-4} = \frac {4}{x^2}\] is this your equation...? @ajl96 ?

OpenStudy (jack1):

\[\Large \frac {6}{x-4} = \frac {4}{x^2} \] --> cross multiply \[\Large \frac {6}{x-4} = \frac {4}{x^2} \] becomes---> \[\Large {6}({x^2}) = {4}({x-4}) \]

OpenStudy (jack1):

\[\Large {6}({x^2}) = {4}({x-4}) \] --> rearrange to form quadratic \[\Large 6x^2 = 4x-16 \] \[\Large 6x^2 - 4x+16 = 0 \] --> use quadratic formula to solve from here